### v2

@@ 2.1 Basic constructions and assumptions
for $p^{k+1}\in\mathbb{L}(X;U)$ computed by taking (single or multiple) Jacobi or Gauss–Seidel steps on the linear equation
(8) $0=T^{(u)}(u^{k+1},x^{k})p^{k+1}+T^{(x)}(u^{k+1},x^{k}).$
Correspondingly, let $S_{w}=S_{p}=S_{u}^{\prime}$ arise from the basic adjoint (4).
If $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable, then, subject to uniformised standard convergence conditions for each of the above-mentioned linear system splitting schemes, Section 2.1Item (ii) holds with $w^{k}=p^{k}$. The proof in [29, Theorem 4.9] is analogous to that of Example 2.5, indeed, (8) has the form (7) with the parameter $x=(u^{k+1},x^{k})$.
Suppose further that
$N_{J^{\prime}}:=\sup\{\|J^{\prime}(S_{u}(x))\|_{U^{*}}\mid x\in\Omega\}<\infty,\quad N_{S_{u}^{\prime}}:=\sup\{\|S_{u}^{\prime}(x)\|_{\mathbb{L}(X;U)}\mid x\in\Omega\}<\infty,$
and that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz. Then Section 2.1Item (iii) holds due to
$\|\widetilde{F^{\prime}}(x^{k})-$ $F^{\prime}(x^{k})\|_{X^{*}}=\|J^{\prime}(u^{k+1})p^{k+1}-J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\|_{X^{*}}$
$\leq\|J^{\prime}(u^{k+1})[p^{k+1}-S_{u}^{\prime}(x^{k})]\|_{X^{*}}+\|[J^{\prime}(u^{k+1})-J^{\prime}(S_{u}(x^{k}))]S_{u}^{\prime}(x^{k})\|_{X^{*}}$
$\leq\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|J^{\prime}(u^{k+1})\|_{U^{*}}+L_{\nabla J}\|S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|u^{k+1}-S_{u}(x^{k})\|_{U}$
$\leq N_{\nabla J}\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}+L_{\nabla J}N_{\nabla S_{u}}\|u^{k+1}-S_{u}(x^{k})\|_{U}.$
Example 2.7 (Reduced adjoint: linear system splitting).
With $U$ and $W_{*}$ Euclidean spaces, and $X$ a normed space, let $S_{u}$ and $T$ be given by (3), and define
@@
for $w^{k+1}\in W$ computed by taking (single or multiple) Jacobi or Gauss–Seidel steps on the linear equation
$w^{k+1}T^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0.$
Correspondingly, let $S_{w}$ arise from the reduced adjoint (5).
If $T|U\times\Omega$, $S_{u}|\Omega$, and $J$ are Lipschitz-continuously differentiable, then Section 2.1Item (ii) holds subject to uniformised standard convergence conditions for each of the above-mentioned linear system splitting schemes. Although not proved in [29], this claim follows similarly to Example 2.6, as both involve linear systems of similar form, with $J^{\prime}$ in the reduced adjoint taking the place of $T^{(x)}$ in the basic adjoint.
Suppose then that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(x)};u}$-Lipshitz for all $x\in\Omega$ with both
$M_{T^{(x)}}:=\sup\{\|T^{(x)}(u,x)\|\mid u\in U,x\in\Omega\}<\infty\enskip\text{and}\enskip N_{S_{w}}:=\sup\{\|S_{w}(x)\|\mid x\in\Omega\}<\infty.$
@@
Then the differential transformation Section 2.1Item (iii) holds due to
$\|\widetilde{F^{\prime}}(x^{k})$ $-F^{\prime}(x^{k})\|_{X^{*}}=\|w^{k+1}T^{(x)}(u^{k+1},x^{k})-w_{x^{k}}T^{(x)}(S_{u}(x^{k}),x^{k})\|_{X^{*}}$
$=\|[w^{k+1}-w_{x^{k}}]T^{(x)}(u^{k+1},x^{k})-w_{x^{k}}[T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})]\|_{X^{*}}$
$\leq\|T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\|w^{k+1}-w_{x^{k}}\|_{W}$
$+\|w_{x^{k}}\|_{W}\|T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}$
$\leq N_{S_{w}}L_{T^{(x)};u}\|u^{k+1}-S_{u}(x^{k})\|_{U}+M_{T^{(x)}}\|w^{k+1}-S_{w}(x^{k})\|_{W}.$
2.2 An error estimate
We now study the error of approximating of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$, for all $k\in\mathbb{N}$. To this end, we first develop a technical estimate for sequences of real numbers satisfying certain recursive inequalities. These abstract estimates will later be instantiated using the inner and adjoint tracking conditions introduced earlier. The main results are core estimates that allow us to isolate the contributions of initialisation and update errors, and thereby quantify the impact of inexact inner and adjoint solutions over multiple iterations on the differential approximations.

### v3

@@ 3.2 Adjoint algorithms and the adjoint tracking inequality; the differential transformation condition
for $p^{k+1}\in\mathbb{L}(X;U)$ computed by taking (single or multiple) admissible splitting steps (see Sections˜3.1, 3.8 and 3.7) on the linear equation
(18) $0=T^{(u)}(u^{k+1},x^{k})p+T^{(x)}(u^{k+1},x^{k})$
with unknown $p=p^{k+1}$. Correspondingly, let $S_{w}=S_{p}=S_{u}^{\prime}$ arise from the basic adjoint (5). The next theorem shows that the scheme satisfies Section˜2.1 in case of single steps; multiple steps follow immediately with larger $\kappa_{w}$ by chaining Section˜2.1Item˜(ii).
Theorem 3.9.
Suppose that $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable for an $\Omega\subset X$, and that the adjoint admissible splitting Section˜3.1 holds in $U\times\Omega$ for the operators $A_{(u,x)}=T^{(u)}(u,x)$. Suppose, moreover, that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ and $T^{(u)}(\,\boldsymbol{\cdot}\,,x)$ are, respectively, $L_{T^{(u)};u}$- and $L_{T^{(x)};u}$-Lipschitz for all $x\in\Omega$; that $S_{u}^{\prime}$ is $L_{S_{u}^{\prime}}$-Lipschitz in $\Omega$; and that
$N_{S_{u}^{\prime}}:=\sup\{\|S_{u}^{\prime}(x)\|_{\mathbb{L}(X;U)}\mid x\in\Omega\}<\infty.$
Let
$p^{k+1}:=N^{-1}_{(u^{k+1},x^{k})}(T^{(x)}(u^{k+1},x^{k})-M_{(u^{k+1},x^{k})}p^{k}).$
Then Section˜2.1Item˜(ii) holds in $\Omega$ for $w^{k}=p^{k}$ with $\kappa_{w}=\zeta^{-1}$, $\mu_{u}=\zeta^{-1}\gamma^{-1}_{N}(L_{T^{(u)};u}N_{S_{u}^{\prime}}+L_{T^{(x)};u})$, and $\pi_{w}=L_{S_{u}^{\prime}}$.
Suppose further that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz and
$N_{J^{\prime}}:=\sup\{\|J^{\prime}(S_{u}(x))\|_{U^{*}}\mid x\in\Omega\}<\infty.$
Then Section˜2.1Item˜(iii) holds for $\widetilde{F^{\prime}}(x^{k}):=J^{\prime}(u^{k+1})p^{k+1}$ with $\alpha_{u}=L_{J^{\prime}}N_{S_{u}^{\prime}}$ and $\alpha_{w}=N_{J^{\prime}}$.
Proof 3.10.
Section˜2.1Item˜(ii) is proved in [30, Theorem 4.9]. That theorem is written entirely in Hilbert spaces, but the proof does not require the setting. Instead of reproducing the proof here, we provide the analogous proof for reduced adjoints in the next Theorem˜3.11. Section˜2.1Item˜(iii) follows from
$\begin{split}\|\widetilde{F^{\prime}}(x^{k})-&F^{\prime}(x^{k})\|_{X^{*}}=\|J^{\prime}(u^{k+1})p^{k+1}-J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\|_{X^{*}}\\ &\leq\|J^{\prime}(u^{k+1})[p^{k+1}-S_{u}^{\prime}(x^{k})]\|_{X^{*}}+\|[J^{\prime}(u^{k+1})-J^{\prime}(S_{u}(x^{k}))]S_{u}^{\prime}(x^{k})\|_{X^{*}}\\ &\leq\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|J^{\prime}(u^{k+1})\|_{U^{*}}+L_{J^{\prime}}\|S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|u^{k+1}-S_{u}(x^{k})\|_{U}\\ &\leq N_{J^{\prime}}\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}+L_{J^{\prime}}N_{S_{u}^{\prime}}\|u^{k+1}-S_{u}(x^{k})\|_{U}.\end{split}$
Reduced adjoint: linear system splitting
With $U$, $W_{*}$, and $X$ normed spaces, let $S_{u}$ and $T$ be given by (1), and define
@@
for $w^{k+1}\in W$ computed by taking (single or multiple) admissible splitting steps (see Sections˜3.1, 3.8 and 3.7) on the linear equation
$wT^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0$
with unknown $w=w^{k+1}$. Correspondingly, let $S_{w}$ arise from the reduced adjoint (6). The next theorem shows that the scheme satisfies Section˜2.1 in case of single steps; multiple steps follow immediately with large $\kappa_{w}$ by chaining Section˜2.1Item˜(ii).
Theorem 3.11.
Suppose that $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable for an $\Omega\subset X$, and that the adjoint admissible splitting Section˜3.1 holds in $U\times\Omega$ for the linear operators $A_{(u,x)}:w\mapsto wT^{(u)}(u,x)$. Suppose, moreover, that $T^{(u)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(u)};u}$-Lipschitz for all $x\in\Omega$; that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz; that $S_{w}$ is $L_{S_{w}}$-Lipschitz in $\Omega$; and that
$N_{S_{w}}:=\sup\{\|S_{w}(x)\|\mid x\in\Omega\}<\infty.$
Let
$w^{k+1}:=-N^{-1}_{(u^{k+1},x^{k})}(J^{\prime}(u^{k+1})+M_{(u^{k+1},x^{k})}w^{k}).$
Then Section˜2.1Item˜(ii) holds in $\Omega$ with $\kappa_{w}=\zeta^{-1}$, $\mu_{u}=\zeta^{-1}\gamma^{-1}_{N}(L_{T^{(u)};u}N_{S_{w}}+L_{J})$, and $\pi_{w}=L_{S_{w}}$.
Suppose further that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(x)};u}$-Lipshitz for all $x\in\Omega$ and that
$M_{T^{(x)}}:=\sup\{\|T^{(x)}(u,x)\|\mid u\in U,x\in\Omega\}<\infty$
@@
Then the differential transformation Section˜2.1Item˜(iii) holds for $\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})$ with $\alpha_{u}=N_{S_{w}}L_{T^{(x)};u}$ and $\alpha_{w}=M_{T^{(x)}}$.
Proof 3.12.
To prove Section˜2.1Item˜(ii), for brevity, let $v=(u^{k+1},x^{k})$, $A_{v}w=wT^{(u)}(v)$, and $b_{v}=-J^{\prime}(u^{k+1})$. Likewise let $\bar{v}=(S_{u}(x^{k}),x^{k})$, $A_{\bar{v}}w=wT^{(u)}(\bar{v})$, and $b_{\bar{v}}=-J^{\prime}(S_{u}(x^{k}))$. Assuming that $x^{k}\in\Omega$, we proceed as in [30, Lemma 4.1]: Observing that $A_{\bar{v}}S_{w}(x^{k})=b_{\bar{v}}$ and $A_{v}S_{w}(x^{k})=(N_{v}+M_{v})S_{w}(x^{k})$, we rearrange
$\begin{split}w^{k+1}-S_{w}(x^{k})&=N^{-1}_{v}(b_{v}-M_{v}w^{k})-S_{w}(x^{k})\\ &=N^{-1}_{v}(b_{v}-b_{\bar{v}})-N^{-1}_{v}(A_{v}-A_{\bar{v}})S_{w}(x^{k})-N^{-1}_{v}M_{v}(w^{k}-S_{w}(x^{k})).\end{split}$
From this it follows, using our assumptions, that
$\begin{split}\zeta^{-1}\|w^{k+1}-S_{w}(x^{k})\|&\leq\zeta^{-1}\gamma^{-1}_{N}(L_{J^{\prime}}+L_{T^{(u)};u}\|S_{w}(x^{k})\|)\|v-\bar{v}\|+\|w^{k}-S_{w}(x^{k})\|\\ &\leq\zeta^{-1}\gamma^{-1}_{N}(L_{J^{\prime}}+L_{T^{(u)};u}N_{S_{w}})\|u^{k+1}-S_{u}(x^{k})\|\\ &+\|w^{k}-S_{w}(x^{k-1})\|+L_{S_{w}}\|x^{k}-x^{k-1}\|.\end{split}$
Section˜2.1Item˜(iii) follows from
$\begin{split}\|\widetilde{F^{\prime}}(x^{k})&-F^{\prime}(x^{k})\|_{X^{*}}=\|w^{k+1}T^{(x)}(u^{k+1},x^{k})-S_{w}(x^{k})T^{(x)}(S_{u}(x^{k}),x^{k})\|_{X^{*}}\\ &=\|[w^{k+1}-S_{w}(x^{k})]T^{(x)}(u^{k+1},x^{k})-S_{w}(x^{k})[T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})]\|_{X^{*}}\\ &\leq\|T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\|w^{k+1}-S_{w}(x^{k})\|_{W}\\ &+\|S_{w}(x^{k})\|_{W}\|T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\\ &\leq N_{S_{w}}L_{T^{(x)};u}\|u^{k+1}-S_{u}(x^{k})\|_{U}+M_{T^{(x)}}\|w^{k+1}-S_{w}(x^{k})\|_{W}.\end{split}$
4 Nonconvex forward-backward type methods with inexact updates
We now need to prove the convergence of outer methods for the overall problem (3), given estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ by inner and adjoint methods, the latter two satisfying the tracking theory of Section˜2. In this section, we do this through a convergence theory for general inexact forward backward-type methods in a normed space $X$. Our treatment encompasses primal-dual methods, seen as forward-backward methods with respect to appropriate operator-relative (semi-)norms, discussed in the previous section. We introduce such methods in Section˜4.1. Then in Section˜4.2 we introduce abstract growth conditions, which we will use in Sections˜4.3, 4.4, 4.5 and 4.6 to prove various forms of convergence.
Afterwards, in Section˜6, we will verify the growth inequalities for forward-backward and primal-dual algorithms that use the tracking theory of Section˜2 for (single-loop) updates of an inner problem.
4.1 General inexact forward-backward type methods
For proper $F,G:X\to\overline{\mathbb{R}}$, consider the problem
$\min_{x\in X}F(x)+G(x).$
In this subsection, and in the examples of Section˜6, $G$ will be convex and lower semicontinuous, and $F$ Fréchet differentiable, but the general theory of Sections˜4.2, 4.3, 4.4, 4.5 and 4.6 will make no such assumption.
For an initial $x^{0}$, if $X$ is Hilbert, the iterates $\{x^{k}\}_{k=1}^{\infty}$ of the basic inexact forward-backward method are generated for some step length parameter $\tau>0$ and an estimate $\widetilde{\nabla F}(x^{k})$ of $\nabla F(x^{k})$ (not necessarily the one from Section˜2) by
In implicit form the method reads
(20) $\tau^{-1}(x^{k+1}-x^{k})\in\widetilde{\nabla F}(x^{k})+\partial G(x^{k+1}).$
(21) $0\in H(x):=F^{\prime}(x)+\partial G(x)+\Xi x$
with an implicit method of the rough form
[…]
Algorithms of the form (22) with an exact inclusion for $\tilde{\partial}_{k+1}$, cover many common splitting algorithms, such as Douglas–Rachford splitting (DRS) and the primal-dual proximal splitting (PDPS) of [4]; see [7, 32]. As we will see in the following examples, with an inexact inclusion, besides the inexact gradients of Section˜2, the approach also covers inexact proximal maps and mismatched adjoints [19] in primal-dual methods. Inexact proximal maps were used, e.g., in [34] for point source localisation in measure spaces.
Example 4.1 (Primal-dual proximal splitting).
(23) $\min_{z\in Z}\penalty 10000\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\penalty 10000\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y).$
If $f$ is convex, subject to the standard condition on the existence of $x_{0}\in\operatorname{int}\operatorname{dom}[h\circ K]\cap\operatorname{dom}[f+g]\neq\emptyset$ with $Kx_{0}\in\operatorname{int}\operatorname{dom}h$, [^1] Several relaxations are possible, include using the relative interior, or the formulas of [2]. the Fenchel–Rockafellar theorem [7, Theorem 5.11] gives rise to the necessary and sufficient first-order primal-dual optimality conditions of the form (21),
where
$F(z,y)=f(z),\quad G(z,y)=g(z)+h_{*}(y),\quad\text{and}\quad\Xi=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$.
If $f$ is nonconvex, the necessity can be shown through, e.g., Mordukhovich subdifferentials, and their compatibility with both convex subdifferentials and Fréchet derivatives; see, e.g., [7].
Pick step length parameters $\tau,\sigma>0$. With inexact gradients for $f$, the PDPS in Hilbert spaces reads
When $f=j\circ S_{u}$ for $S_{u}$ a PDE solution operator, and we compute $\widetilde{\nabla f}$ following Theorems˜3.5 and 3.11, (24) becomes the algorithm presented in [18].
To extend (24) to general normed spaces, we write it in $X=Z\times Y$ in the implicit form (22) as
(25a) $0\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}+M(x^{k+1}-x^{k})\quad\text{for}$
(25b) $\widetilde{F^{\prime}}(z^{k},y^{k}):=\begin{pmatrix}\widetilde{f^{\prime}}(z^{k})\\ 0\end{pmatrix}\quad\text{and}\quad M:=\begin{pmatrix}\tau^{-1}M_{z},&-K^{*}\\ -K&\sigma^{-1}M_{y}\end{pmatrix}$
4.2 Inexact growth inequalities
We now provide several alternative assumptions that give a precise meaning to the approximate inclusion in (22). For this, we first define the Lagrangian gap functional
Example 4.2.
Example 4.3.
For the PDPS of Example˜4.1, with $x=(y,z)$, we obtain the Lagrangian duality gap
Assumption 4.3.
  1. (i)
    (26) $\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{\mathrm{desc},k}.$
  2. (ii)
  3. (iii)
  4. (iv)
    For some $0\leq\tilde{\eta}<\eta$, we have
Remark 4.4.
If $\Omega=X$, convergence will be global. In the examples of Section˜3, $\Omega\neq X$ may arise from $S_{u}$, $G$, or $J$ being only locally Lipschitz continuously differentiable.
We will also need the approximate subdifferentials $\tilde{\partial}_{k+1}$ to become better as the distance between the iterates shrinks, in the sense of
Assumption 4.4.
For $H$ defined in (21), we have
(27) $\langle\tilde{\partial}_{k+1}-H(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}),$
for all $k\in\mathbb{N}$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ for an open neighbourhood $\Omega_{\bar{x}}$ of $\bar{x}$, a positive semi-definite self-adjoint $\breve{\Lambda}\in\mathbb{L}(X;X^{*})$, errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$, and a $\gamma\geq 0$. We recall here that $H(\bar{x})$ is a set, and the notation (27) means that the inequality holds for all elements of this set.
(28) $\langle\tilde{\partial}_{k+1}|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
Let $G$ be convex and $F$ possibly nonconvex but with $L$-Lipschitz Fréchet derivative. To model the basic exact forward-backward splitting in our general framework, we take $\Xi=0$ and $M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial G(x^{k+1})+F^{\prime}(x^{k})$. Then Eq.˜28 with $\breve{\Lambda}=L\operatorname{Id}$, $\gamma=0$, and $\varepsilon_{k}(\bar{x})=0$ is an immediate consequence of the definition of the convex subdifferential for $G$, and of the three-point descent inequality (10) for $F$. The monotonicity (27) likewise follows from the monotonicity of $\partial G$ and a three-point monotonicity property of $\nabla F$; see [7, Chapter 7].
The errors will also need to have a finite sum:
Assumption 4.5.
  1. (a)
    (27) holds, and $\bar{x}\in H^{-1}(0)$; or
  2. (b)
    (28) holds, and $\inf_{x\in\Omega_{\bar{x}}}\mathcal{G}(x;\bar{x})\geq 0.$
[…]
We first show the potentially global convergence of subdifferentials; see Remark˜4.4. When $\Xi=0$, this could be followed by the Kurdyka–Łojasiewicz property to show function value convergence, and, afterwards, either by a growth condition or, in finite dimensions, a finite-length argument based on (30) and [1, proof of Lemma 2.6] to show iterate convergence. As the property can easily be verified only in finite dimensions (for semi-algebraic functions), we prefer a more direct approach.
Theorem 4.6.
If Section˜4.2 holds, then $x^{k}\in\Omega$ and
If, moreover, Section˜4.2 holds, then also $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$
Proof 4.7.
By the implicit algorithm (22), the properties of Fenchel conjugates (e.g., [7, Lemma 5.7]) and $M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)(x^{k+1}-x^{k})$, we have
If $\{x^{j}\}_{j=0}^{N-1}\subset\Omega$, Section˜4.2Item˜(i) thus yields for all $k=0,\ldots,N-1$ that
Summing over all such $k$, and using Section˜4.2Item˜(ii), it follows
From Section˜4.2Item˜(iii), it now follows that $x^{N}\in\Omega$. Since, by the same assumption, $x^{0}\in\Omega$, induction establishes (30) and $x^{k}\in\Omega$ for all $k\in\mathbb{N}$. Using Section˜4.2Item˜(iv) in (33), we, moreover, deduce $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ and $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})\to 0$. Let $c\geq\|M\|_{\mathbb{L}(X;X^{*})}$. By $\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\leq c\|\,\boldsymbol{\cdot}\,\|_{X}^{2}$ and the properties of conjugates (e.g., [7, Lemmas 5.4 and 5.7]),
Thus also $\|\tilde{\partial}_{k+1}\|_{X^{*}}\to 0$. Section˜4.2 proves that $\inf_{x^{*}\in H(x^{k+1})}\|\tilde{\partial}_{k+1}-x^{*}\|_{X^{*}}\to 0$. Hence an application of the triangle inequality establishes $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$.
Remark 4.8 (Forward-backward splitting).
For the (inexact) forward-backward splitting of (20), once we verify the relevant assumptions in Corollary˜6.7, the previous theorem establishes the monotonicity of function values, as well as the convergence of subdifferentials to zero, $\inf_{x^{*}\in\partial G(x^{k+1})}\|F^{\prime}(x^{k+1})+x^{*}\|\to 0$.
Remark 4.9 (Bregman divergences).
The argument of Theorem˜4.6 extends to algorithms where the $M$-seminorm for $M\in\mathbb{L}(X;X^{*})$ is replaced by a Bregman divergence $B_{M}$ generated by some convex $M:X\to\overline{\mathbb{R}}$. Recalling the definition
in the algorithm (22), we would impose $\tilde{\partial}_{k+1}=\omega^{k+1}-\omega^{k}\in\partial_{2}B_{M}^{\omega^{k}}(x^{k},x^{k+1})$ for a given $\omega^{k}\in\partial M(x^{k})$ and some $\omega^{k+1}\in\partial M(x^{k+1})$. Using the Fenchel–Young identity, we could then replace (31) by
Thus repeating the arguments of the theorem would establish both $B_{M}^{\omega^{k}}(x^{k},x^{k+1})\to 0$ as well as $B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})=B_{M^{*}}^{x^{k+1}}(\omega^{k}+\tilde{\partial}_{k+1},\omega^{k})\to 0$. A variant of Section˜4.2 could then establish a form of convergence for $H(x^{k+1})$.
4.4 Non-escape, quasi-Féjer monotonicity, linear convergence
The next lemma is essential for all our strong convergence results. The proof is standard; see, e.g., [7, Chapter 15] for the case $\varepsilon_{k}(\bar{x})=0$ and $\Xi=0$. Observe that (35) with the triangle inequality may be used to again prove Section˜2.1Item˜(i) for multilevel methods.